# Missing Money and Market-Based Adequacy in Deeply Decarbonized Power Systems with Long-Duration Energy Storage

### v1

@@ IV-D CM impacts on prices and LDES net revenues
Introduction of CM affects energy prices and LDES revenues. Table III contains consumer cost and price volatility metrics. Tighter emission limits raise energy costs. Due to price caps and payments through CMs, energy prices in E+CM are lower by 15–30 \$/MWh compared to EOM, but capacity procurement costs, passed to consumers, offset these savings. Thus, consumer costs are similar in EOM and E+CM, indicating no extra financial burden for the CM. [^10] Administrative and transaction costs, which could marginally impact overall costs, are not considered in this analysis. CMs also lower energy price volatility, though this effect diminishes with stricter emission limits.
TABLE III: Price and Cost Metrics for All Runs
@@ IV-E Sensitivity to capacity credit estimation inaccuracies
We assess the impact of capacity credit estimation errors on market performance through a sensitivity analysis, adjusting the capacity credit of storage technologies to account for the dependence of CC on the resource mix and potential inaccuracies of the accreditation methodologies. For each simulation, we use the adjusted credits to estimate a new capacity target. We keep the net-CONE for the reference technology the same across all runs. Fig. 6 shows the change in the average energy price, capacity price, social welfare, and EUE. Under credit underestimation, the investment signal provided by the capacity market to storage weakens, leading to lower investment and tighter conditions in the energy market. As a result of lower capacity targets, capacity market prices reduce, and energy prices and EUE increase. The changes are opposite under credit overestimation. We also find an asymmetric effect on welfare, with underestimation causing larger losses. The SW marginal increases with overestimated credits for 1 and 0.1 gCO2/kWh cases. In the former case, this is due to misaligned credit (per Section IV-B), and in the latter, due to the mix that better negates the impact of the price cap on storage dispatch.
@@ V Discussion and conclusions
The analysis is performed for a case study with renewable and load profiles that resemble those of Great Britain. Simulations consider a range of CO2 emission targets that yield varying levels of LDES integration. We find that price cap distortions result in a substantial missing money problem for LDES (approximately 15% to 50% of LDES costs are not recovered when price caps are retroactively applied to an optimal capacity mix). The magnitude of the problem reduces under more decarbonized systems as storage can capture higher spreads due to low charging costs during periods of abundant renewable supply and high discharging costs during periods with expensive carbon-emitting marginal suppliers.
@@
We demonstrate that well-calibrated CMs can provide near-efficient investment signals, resulting in welfare outcomes comparable to those of an energy-only market without price caps. These results expand the work of Wang et al. [21], by showing that efficient outcomes are possible even when demand is flexible, storage duration is an investment decision, and ES makes up a significant share of installed capacity. In cases of deeper decarbonization, we find that the capacity market is less effective in restoring system security of supply, as measured by expected unserved energy, to the levels that would prevail without price caps. Two factors explain this result. First, the capacity credits required for cost recovery are not always aligned with those estimated through a marginal accreditation approach, as the latter is highly sensitive to assumptions about storage dispatch. Second, in the presence of storage, energy price caps distort market outcomes beyond periods of scarcity. By capping discharging revenues, the price cap undermines incentives for charging in non-scarcity hours. This suggests that conventional methods that only consider the difference between PC and VOLL may be poorly suited to parameterize capacity markets in deeply decarbonized systems.
@@
Our work provides a series of timely regulatory insights. While CMs can, in principle, remain an effective tool for ensuring adequacy, they are complex and sensitive to parameter values such as capacity credits. Nonetheless, CMs effectiveness can remain high in the medium term, provided that accreditation methods are transparent and predictable, giving investors greater certainty. In the longer term, the definition of missing money itself may need to be reevaluated, as it should account for how price caps distort storage revenues across both scarcity and non-scarcity periods. Given that the ability of CMs to reduce price and revenue volatility declines with deeper decarbonisation, complementary measures, such as risk-hedging government-backed contracts, are worth further investigation.
Future work on business models for LDES can extend the framework presented in this article by modeling risk-averse investors, considering relevant policy mechanisms and novel electricity market products that might be a good fit for LDES, and limiting the horizon within which LDES has foresight into future energy prices.

### v2

@@ IV-D CM impacts on prices and LDES net revenues
Introducing a capacity market affects energy prices and LDES revenues. Table V contains consumer cost and price volatility metrics. Tighter emission limits raise energy costs. Due to price caps and payments through CMs, energy prices in E+CM are lower by 15–30 \$/MWh compared to EOM VOLL, but capacity procurement costs, passed to consumers, offset these savings. Thus, consumer costs are similar in EOM VOLL and E+CM, indicating no extra financial burden for the CM. [^11] Administrative and transaction costs, which could marginally impact overall costs, are not considered in this analysis. CMs also lower energy price volatility, though this effect diminishes with stricter emission limits.
TABLE V: Price and Cost Metrics for All Runs
@@ IV-E Sensitivity to capacity credit estimation inaccuracies
We assess the impact of capacity credit estimation errors on market performance through a sensitivity analysis, adjusting the capacity credit of storage technologies to account for the dependence of CC on the resource mix and potential inaccuracies of the accreditation methodologies. For each simulation, we use the adjusted credits to estimate a new capacity target. We keep the net-CONE for the reference generator the same across all sensitivities. Fig. 7 shows the change in the average energy price, capacity price, social welfare, and EUE. Under credit underestimation, the investment signal provided by the capacity market to storage weakens, leading to lower investment and tighter conditions in the energy market. As a result of lower capacity targets, capacity market prices reduce, and energy prices and EUE increase. The changes are opposite under credit overestimation. We also find an asymmetric effect on welfare, with underestimation causing larger losses. The SW marginal increases with overestimated credits for 1 and 0.1 gCO2/kWh cases. In the former case, this is due to weaker relationship between missing money and EUE contributions (per Section IV-B), and in the latter, due to the mix that better negates the impact of the price cap on storage dispatch.
@@ V Discussion and conclusions
The analysis is performed for a case study with renewable and load profiles that resemble those of Great Britain. Simulations consider a range of CO2 emission targets that yield varying levels of LDES integration. We find that price cap distortions result in a substantial missing money problem for LDES (approximately 15% to 50% of LDES costs are not recovered when a baseline mix identified by maximising SW is dispatched with energy price caps). The magnitude of the problem reduces under more decarbonized systems, as storage can capture higher spreads due to low charging costs during periods of abundant renewable supply and high discharging costs during periods with expensive carbon-emitting marginal suppliers.
@@
We demonstrate that well-calibrated CMs can provide near-efficient investment signals, resulting in welfare outcomes comparable to those of an energy-only market without price caps. These results expand the work of Wang et al. [wang_crediting_2022], by showing that efficient outcomes are possible even when demand is flexible, storage duration is an investment decision, and ES makes up a significant share of installed capacity. In cases of deeper decarbonization, we find that the capacity market is less effective in restoring system security of supply, as measured by expected unserved energy, to the levels that would prevail without price caps. Two factors explain this result. First, the relationship between missing money and marginal contributions to adequacy varies per technology, when demand is price-elastic, and storage is charging-constrained. Second, in the presence of storage, energy price caps distort market outcomes beyond periods of scarcity. By capping discharging revenues, the price cap undermines incentives for charging in non-scarcity hours. This suggests that conventional methods that only consider the difference between price cap and VOLL may be poorly suited to parameterize capacity markets in deeply decarbonized systems.
@@
Our work provides a series of timely regulatory insights. While CMs can, in principle, remain an effective tool for ensuring adequacy, they are complex and sensitive to parameter values such as capacity credits. Nonetheless, CMs effectiveness can remain high in the medium term, provided that accreditation methods are transparent and predictable, giving investors greater certainty. In the longer term, capacity market parameterization methodologies may need to be refined to reflect how price caps, in systems with high storage penetration, affect revenue formation and reliability contribution not only during scarcity hours but also outside scarcity periods. Given that the ability of CMs to reduce price and revenue volatility declines with deeper decarbonisation, complementary measures, such as risk-hedging government-backed contracts, are worth further investigation.
Future work on business models for LDES can extend the framework presented in this article by modeling risk-averse investors, considering relevant policy mechanisms and novel electricity market products that might be a good fit for LDES, and limiting the horizon within which LDES has foresight into future energy prices. Performance penalties are another key design element of CRMs that merits deeper investigation, as their enforcement is complicated by the presence of charging-constrained conditions. Furthermore, the framework can be extended to include network constraints and an explicit representation of the system operator. This could be implemented through zonal or nodal formulations that account for transmission limits and locational marginal pricing, with the system operator acting as an additional agent ensuring network feasibility of the energy schedule. Such extensions would allow the assessment of spatial adequacy and locational incentives under alternative market structures. Finally, an important avenue for future research is the development of more fundamental mathematical conditions under which marginal EFC-based accreditation restores long-run cost recovery and adequacy in systems with price-elastic demand and charging-constrained storage.